OpenStudy (anonymous):

You were billed $17.60 for water usage at a rate of $0.55 per gallon. Three choices are correct ways to calculate the the number of gallons, g, of water that you used. Which choice is NOT? a) 17.6 = 0.55 • g b) 17.60/55 = g c) 0.55g = 17.6 d)0.55/g = 17.6

OpenStudy (anonymous):

Looks like a typo in b). Should it be 17.60/0.55 = g ?

OpenStudy (texaschic101):

basically the correct way would be to divide 17.60 by 0.55 to calculate the number of gallons used. (a) 17.6 = 0.55g ....to solve this, you would divide both sides by 0.55 1.7.6/0.55 = g (this is correct) (b) 17.60/0.55 = g (this is correct as well) (c) 0.55g = 17.6...you would divide both sides by 0.55 g = 17.6/0.55 (this is correct) (d) 0.55/g = 17.5.....to solve this, you would multiply both sides by g 0.55 = 17.5/g (incorrect)

OpenStudy (anonymous):

ahh okay, thank you